We find a sample of people and we weigh each person. The distribution of their weights is positively skewed with a mean of 157 and a standard deviation of 47. If this distribution is transformed into z-scores, what will be the resulting shape, mean, and standard deviation of the new distribution?

For a population with µ = 75 and σ = 10 find the z-score corresponding to the following raw scores X = 70 X = 77 X = 75
Answer #1

The distribution of their weights is positively skewed with a mean of 157 and a standard deviation of 47.

If this distribution is transformed into z-score, then shape of the new distribution is bell shape.

Mean = 0

Standard deviation = 1

( By using central limit theorem)

For a population with µ = 75 and σ = 10

We have to find z score.

1 -

x = 70

70-75 10 = = -0.5

x = 77

77-75 10 - = 0.2

x = 75

75-75 10 - = 0